# Elucidated Rolling Diffusion Models for Probabilistic Weather Forecasting

### v1

@@ C.3 Stochastic Heun Sampler
In the main text, we introduced our proposed deterministic, first-order sampling algorithm (Algorithm 2). We now extend this into a more sophisticated sampler, detailed in Algorithm 3, by incorporating two key modifications: a second-order Heun correction step and optional sampling stochasticity via a “churn” mechanism. Both components are inspired by EDM’s stochastic sampling algorithm [23], but require specific adaptations for our ERDM to work.
@@
The Heun method enhances sampling accuracy by incorporating a corrector step over the basic Euler step during the denoising of the data. However, a complication arises because when $t_{\rm{next}}^{\prime}\geq 1$, the first snapshot of $\bar{\bm{x}}_{\rm{next}}^{\prime}$ is already fully denoised (i.e., “clean”). Applying the denoiser to such a clean snapshot as part of the Heun step would be inappropriate. To address this, we pad the noisy data by one extra pure-noise snapshot to the end of the sequence, resulting in a window of size $W+1$. The core denoiser, $D_{\theta}$, is then selectively applied on the sub-window of size $W$, depending on whether the first snapshot is already clean or not. That is, if $t_{\rm{next}}^{\prime}\geq 1$, the denoiser would be applied on the snapshots with indices $12,\dots,W+$ 1. Otherwise, it is applied, as usual, to the indices $1,\dots,W$. To formalize this selective application, we define $D_{\theta}^{\mathrm{pad}}({\bm{x}},\bm{\bar{\sigma}}(t))$ such that it applies the actual denoiser on all snapshots with indices $I=\{\lfloor t\rfloor+1,\dots,\lfloor t\rfloor+W\}$ and leaves all other snapshots untouched. This ensures that the denoiser evaluations within the Heun step are confined to snapshots that are indeed noisy.
$\vspace{-4mm}D_{\theta}^{\mathrm{pad}}({\bm{x}},\bm{\bar{\sigma}}(t))_{w}:=\begin{cases}D_{\theta}({\bm{x}}_{I},\bm{\bar{\sigma}}(t)_{I})_{w-\lfloor t\rfloor}&\text{if }w\in I\\ {\bm{x}}_{w}&\text{otherwise}\end{cases}\vspace{-8mm}$ (23)
@@
1:Require: $\hat{{\bm{y}}}_{1:W}^{\mathrm{init}},N,T_{\mathrm{forecast}},S_{\text{churn}}=0,S_{\text{noise}}=1$triangleright Choose $S_{\text{churn}}>0$ for sampling stochasticity
@@
3:$\Delta t\leftarrow 1/N$triangleright Infer step size from given number of steps per snapshot
4:$\Delta t^{\prime}\leftarrow\Delta t/(1-S_{\text{churn}})$  $\triangleright$ Size of larger, initial step, before backtracking (if $S_{\text{churn}}>0$) to $\Delta t$
5:$t_{\rm{cur}}\leftarrow 0;\;\mathcal{S}\leftarrow\emptyset$  $\triangleright$ Initialize global diffusion time and empty generated sequence
6:sample $\bar{\bm{x}}_{\rm{cur}}\sim\mathcal{N}(\hat{{\bm{y}}}_{1:W}^{\mathrm{init}},\bm{\bar{\sigma}}(t_{\rm{cur}})^{2}\mathbf{I}_{W\times D})$triangleright Initialize window with snapshot-dependent rolling noise
7:concatenate $n_{\mathrm{pad}}:=\lfloor\Delta t^{\prime}\rfloor+1$ fully noisy snapshots to the end of $\bar{\bm{x}}_{\rm{cur}}$triangleright Necessary for Heun step
@@
9:    $t_{\rm{next}}^{\prime}\leftarrow t_{\rm{cur}}+\Delta t^{\prime}$triangleright Global diffusion time after denoising, before churn
10:    $t_{\rm{next}}\leftarrow t_{\rm{cur}}+\Delta t$triangleright Global diffusion time after churn ($t_{\rm{next}}\leq t_{\rm{next}}^{\prime}$)
11:    $\bm{\sigma}_{\rm{cur}}\leftarrow\bm{\bar{\sigma}}(t_{\rm{cur}})$triangleright Current noise levels
12:    $\bm{\sigma}_{\rm{next}}^{\prime}\leftarrow\bm{\bar{\sigma}}(t_{\rm{next}}^{\prime})$triangleright Noise levels before churn
13:    $\bm{\sigma}_{\rm{next}}\leftarrow\bm{\bar{\sigma}}(t_{\rm{next}})$triangleright Noise levels at the end of this iteration
@@
15:    $\hat{{\bm{y}}}\leftarrow D_{\theta}^{\mathrm{pad}}(\bar{\bm{x}}_{\rm{cur}},\bm{\sigma}_{\rm{cur}})$triangleright Denoise sequence
16:    $\bm{d}\leftarrow(\bar{\bm{x}}_{\rm{cur}}-\hat{{\bm{y}}})/\bm{\sigma}_{\rm{cur}}$triangleright Evaluate $\mathrm{d}\bm{x}/\mathrm{d}\tau$ at $t_{\rm{cur}}$
17:    $\bar{\bm{x}}_{\rm{next}}^{\prime}\leftarrow\bar{\bm{x}}_{\rm{cur}}+(\bm{\sigma}_{\rm{next}}^{\prime}-\bm{\sigma}_{\rm{cur}})\bm{d}$triangleright Euler step from $t_{\rm{cur}}$ to $t_{\rm{next}}^{\prime}$
@@
19:    $\bm{d}^{\prime}\leftarrow(\bar{\bm{x}}_{\rm{next}}^{\prime}-D_{\theta}^{\mathrm{pad}}(\bar{\bm{x}}_{\rm{next}}^{\prime},\bm{\sigma}_{\rm{next}}^{\prime}))/\bm{\sigma}_{\rm{next}}^{\prime}$triangleright Denoise at $t_{\rm{next}}^{\prime}$
20:    $\bar{\bm{x}}_{\rm{next}}^{\prime}\leftarrow\bar{\bm{x}}_{\rm{cur}}+\frac{1}{2}(\bm{\sigma}_{\rm{next}}^{\prime}-\bm{\sigma}_{\rm{cur}})(\bm{d}+\bm{d}^{\prime})$triangleright $2^{\text{nd}}$ order Heun correction step
@@
22:    sample $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},S_{\text{noise}}^{2}\mathbf{I}_{W\times D})$triangleright Sample churn noise
@@
25:    $n_{\rm{clean}}\leftarrow\lfloor t_{\rm{next}}\rfloor$triangleright Infer finished snapshots. If 00, the following lines are a no-op
@@
28:    $\bar{\bm{x}}_{\rm{cur}}\leftarrow[\bar{\bm{x}}_{\rm{next}},{\bm{x}}_{\mathrm{new}}]$triangleright Concatenate fresh noisy snapshots to futuremost
29:    $t_{\rm{cur}}\leftarrow t_{\rm{next}}-n_{\rm{clean}}$triangleright Re-adjust global diffusion ‘time’ to be in $[0,1)$
@@
For the experiments reported in this work, the stochastic Heun sampler (Algorithm 3) is always used with the Heun correction step. Our ablations in Appendix E demonstrate that our second-order sampler provides better results than a first-order version (Euler-only; Algorithm 2). However, we disable sampling stochasticity by setting the churn rate $S_{\text{churn}}=0$. In our experience, we found that small values of $S_{\text{churn}}\approx 0.1$ can increase the ensemble spread, albeit it had minimal impact on error metrics such as the CRPS. We defer a more comprehensive analysis of stochastic sampling to future work. Our algorithm can, in principle, be used to fully denoise multiple snapshots at once (e.g., output two snapshots per iteration when $N=\frac{1}{2}$). This could be useful in certain modalities or problems where subsequent snapshots are highly correlated, but is left to future work too.
@@ C.4 Temporal noise prior
In our training and sampling algorithms (Alg. 1-3), we use i.i.d. noise to corrupt the first window as well as append new pure-noise snapshots when sliding the window. As mentioned in the main text, it has been shown to be useful to sample temporally correlated noise snapshots. In our experiments, we use the simple “progressive noise model” technique proposed by Ge et al. [15], which we found to improve results on the Navier-Stokes dataset compared to using i.i.d. noise (see Appendix E). This noise prior generates the noise for each snapshot autoregressively by combining a perturbed version of the previous noise with a new i.i.d. noise sample. Mathematically, the noise for the first snapshot is random, $\bm{\epsilon}^{(1)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, while for the following snapshots $k>1$:
@@
where $\alpha$ controls the correlation strength between snapshot noises and $\alpha=0$ corresponds to i.i.d. noise. In our experiments, we use $\alpha=1$ following the recommended practice in [15]. These temporally correlated noise samples replace the i.i.d. draws in lines 6 and 27 of Algorithm 3. For example, line 6 becomes $\bar{\bm{x}}_{\rm{cur}}\leftarrow\hat{{\bm{y}}}_{1:W}^{\mathrm{init}}+\bm{\bar{\sigma}}(t_{\rm{cur}})\bm{\epsilon}_{1:W}$, where $\bm{\epsilon}_{1:W}:=(\bm{\epsilon}^{(1)},\dots,\bm{\epsilon}^{(W)})$. Note that the noise prior is applied during both training and sampling, enabling the model to learn how to deal with temporally correlated noise samples. A more comprehensive study of alternative noise priors in the context of ERDM is left to future work.
@@ D.1 Dataset details
We use the Navier-Stokes dataset with four obstacles introduced by Otness et al. [38]. Its license is CC BY 4.0. We also use the ERA5 dataset provided by Weatherbench-2 [43]. Its license is CC BY 4.0.
@@
We use the $1.5^{\circ}$ resolution of the data (a $240\times 121$ grid) provided by [43], with 69 prognostic (input and output) variables: Temperature (t), geopotential (z), specific humidity (q), and the u and v components of wind (u,v) over 13 pressure levels as well as the four surface variables 2m temperature (2t), mean sea level pressure (mslp), and u and v components of winds at 10m (10u, 10v). Following Weatherbench-2, the 13 levels are $\{50,100,150,200,250,300,400,500,600,700,850,925,1000\}$. Numbers after any variable abbreviation refer to the level in hPa. For example, z500 refers to geopotential at 500 hPa. Besides these prognostic variables, we also use six static maps as additional inputs: A land-sea mask, soil type, orography, and $\sin(lat),\cos(lat)\cos(lon),\cos(lat)\sin(lon)$, where $lat,lon$ are the latitude and longitude of each location. We train our models on $12$-hourly data, as in [42], from 1979 to 2020 and evaluate the models based on 64 initial conditions at 00/12 UTC, evenly spaced in 2021.
@@ D.2 Implementation details and hyperparameters
For all experiments and baselines that we train, we use the same basic ADM U-Net [12] architecture, only modified where necessary depending on the method. We always use $4$ down- and up-sampling blocks, with channel multipliers $1,2,3,4$. The base channel dimension is $64$ for Navier-Stokes and $256$ for ERA5 experiments. We use attention layers at the two lowest spatial resolutions. We use a dropout rate of 0.15 for Navier-Stokes, but disable it for ERA5. Other architectural hyperparameters are kept to the defaults from EDM. This results in a parameter count of $29.5$ million for Navier-Stokes, and $517$ million for ERA5. For the models operating on sequence, ERDM and EDM with $W>1$, we integrate causal temporal attention layers into the 2D U-Net before each down- and up-sampling block. This increases the respective parameter counts slightly to $31.5$ and $537$ million for Navier-Stokes and ERA5, respectively. For ERA5, we use circularly padded convolutions along the longitude dimension (and zero-padding for the latitude dimension) to better respect the spherical structure of the Earth.
@@
We train each model for 300 epochs with a global effective batch size of 32, leading to almost $57,000$ gradient steps. The effective batch size is kept the same by balancing per-GPU batch size and gradient accumulation steps appropriately over different training runs. We use a cosine learning rate schedule with a linear warmup period of 10 epochs and a peak learning rate of $6\times 10^{-4}$. We use the AdamW optimizer with a weight decay of $10^{-4}$. During evaluation, we use an exponential moving average (EMA)–based on an EMA decay rate of $0.995$–version of the model. Because the input spatial grid, $221\times 42$, after four halvings does not result in an even number–a problem for the residual connections in the upsampling blocks–we bilinearly interpolate it to a $256\times 64$ grid at the input layer (and back before the final convolution layer).

### v3

@@ C.3 Stochastic Heun Sampler
In the main text, we introduced our proposed deterministic, first-order sampling algorithm (Algorithm˜2). We now extend this into a more sophisticated sampler, detailed in Algorithm˜3, by incorporating two key modifications: a second-order Heun correction step and optional sampling stochasticity via a “churn” mechanism. Both components are inspired by EDM’s stochastic sampling algorithm [karras2022edm], but require specific adaptations for our ERDM to work.
@@
The Heun method enhances sampling accuracy by incorporating a corrector step over the basic Euler step during the denoising of the data. However, a complication arises because when $t_{\rm{next}}^{\prime}\geq 1$, the first snapshot of $\bar{\bm{x}}_{\rm{next}}^{\prime}$ is already fully denoised (i.e., “clean”). Applying the denoiser to such a clean snapshot as part of the Heun step would be inappropriate. To address this, we pad the noisy data by one extra pure-noise snapshot to the end of the sequence, resulting in a window of size $W+1$. The core denoiser, $D_{\theta}$, is then selectively applied on the sub-window of size $W$, depending on whether the first snapshot is already clean or not. That is, if $t_{\rm{next}}^{\prime}\geq 1$, the denoiser would be applied on the snapshots with indices $2,\dots,W+1$. Otherwise, it is applied, as usual, to the indices $1,\dots,W$. To formalize this selective application, we define $D_{\theta}^{\mathrm{pad}}({\bm{x}},\bm{\bar{\sigma}}(t))$ such that it applies the actual denoiser on all snapshots with indices $I=\{\lfloor t\rfloor+1,\dots,\lfloor t\rfloor+W\}$ and leaves all other snapshots untouched. This ensures that the denoiser evaluations within the Heun step are confined to snapshots that are indeed noisy.
$\vskip-11.38109ptD_{\theta}^{\mathrm{pad}}({\bm{x}},\bm{\bar{\sigma}}(t))_{w}:=\begin{cases}D_{\theta}({\bm{x}}_{I},\bm{\bar{\sigma}}(t)_{I})_{w-\lfloor t\rfloor}&\text{if }w\in I\\ {\bm{x}}_{w}&\text{otherwise}\end{cases}\vskip-22.76219pt$ (23)
@@
1:Require: $\hat{{\bm{y}}}_{1:W}^{\mathrm{init}},N,T_{\mathrm{forecast}},S_{\text{churn}}=0,S_{\text{noise}}=1$ $\triangleright$ Choose $S_{\text{churn}}>0$ for sampling stochasticity
@@
[…]
6:sample $\bar{\bm{x}}_{\rm{cur}}\sim\mathcal{N}(\hat{{\bm{y}}}_{1:W}^{\mathrm{init}},\bm{\bar{\sigma}}(t_{\rm{cur}})^{2}\mathbf{I}_{W\times D})$ $\triangleright$ Initialize window with snapshot-dependent rolling noise
7:concatenate $n_{\mathrm{pad}}:=\lfloor\Delta t^{\prime}\rfloor+1$ fully noisy snapshots to the end of $\bar{\bm{x}}_{\rm{cur}}$ $\triangleright$ Necessary for Heun step
@@
9:  $t_{\rm{next}}^{\prime}\leftarrow t_{\rm{cur}}+\Delta t^{\prime}$ $\triangleright$ Global diffusion time after denoising, before churn
10:  $t_{\rm{next}}\leftarrow t_{\rm{cur}}+\Delta t$ $\triangleright$ Global diffusion time after churn ($t_{\rm{next}}\leq t_{\rm{next}}^{\prime}$)
11:  $\bm{\sigma}_{\rm{cur}}\leftarrow\bm{\bar{\sigma}}(t_{\rm{cur}})$ $\triangleright$ Current noise levels
12:  $\bm{\sigma}_{\rm{next}}^{\prime}\leftarrow\bm{\bar{\sigma}}(t_{\rm{next}}^{\prime})$ $\triangleright$ Noise levels before churn
13:  $\bm{\sigma}_{\rm{next}}\leftarrow\bm{\bar{\sigma}}(t_{\rm{next}})$ $\triangleright$ Noise levels at the end of this iteration
@@
15:  $\hat{{\bm{y}}}\leftarrow D_{\theta}^{\mathrm{pad}}(\bar{\bm{x}}_{\rm{cur}},\bm{\sigma}_{\rm{cur}})$ $\triangleright$ Denoise sequence
16:  $\bm{d}\leftarrow(\bar{\bm{x}}_{\rm{cur}}-\hat{{\bm{y}}})/\bm{\sigma}_{\rm{cur}}$ $\triangleright$ Evaluate $\mathrm{d}\bm{x}/\mathrm{d}\tau$ at $t_{\rm{cur}}$
17:  $\bar{\bm{x}}_{\rm{next}}^{\prime}\leftarrow\bar{\bm{x}}_{\rm{cur}}+(\bm{\sigma}_{\rm{next}}^{\prime}-\bm{\sigma}_{\rm{cur}})\bm{d}$ $\triangleright$ Euler step from $t_{\rm{cur}}$ to $t_{\rm{next}}^{\prime}$
@@
19:  $\bm{d}^{\prime}\leftarrow(\bar{\bm{x}}_{\rm{next}}^{\prime}-D_{\theta}^{\mathrm{pad}}(\bar{\bm{x}}_{\rm{next}}^{\prime},\bm{\sigma}_{\rm{next}}^{\prime}))/\bm{\sigma}_{\rm{next}}^{\prime}$ $\triangleright$ Denoise at $t_{\rm{next}}^{\prime}$
20:  $\bar{\bm{x}}_{\rm{next}}^{\prime}\leftarrow\bar{\bm{x}}_{\rm{cur}}+\frac{1}{2}(\bm{\sigma}_{\rm{next}}^{\prime}-\bm{\sigma}_{\rm{cur}})(\bm{d}+\bm{d}^{\prime})$ $\triangleright$ $2^{\text{nd}}$ order Heun correction step
@@
22:  sample $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},S_{\text{noise}}^{2}\mathbf{I}_{W\times D})$ $\triangleright$ Sample churn noise
@@
25:  $n_{\rm{clean}}\leftarrow\lfloor t_{\rm{next}}\rfloor$ $\triangleright$ Infer finished snapshots. If 0, the following lines are a no-op
@@
28:  $\bar{\bm{x}}_{\rm{cur}}\leftarrow[\bar{\bm{x}}_{\rm{next}},{\bm{x}}_{\mathrm{new}}]$ $\triangleright$ Concatenate fresh noisy snapshots to futuremost
29:  $t_{\rm{cur}}\leftarrow t_{\rm{next}}-n_{\rm{clean}}$ $\triangleright$ Re-adjust global diffusion ‘time’ to be in $[0,1)$
@@
For the experiments reported in this work, the stochastic Heun sampler (Algorithm˜3) is always used with the Heun correction step. Our ablations in Appendix˜E demonstrate that our second-order sampler provides better results than a first-order version (Euler-only; Algorithm˜2). However, we disable sampling stochasticity by setting the churn rate $S_{\text{churn}}=0$. In our experience, we found that small values of $S_{\text{churn}}\approx 0.1$ can increase the ensemble spread, albeit it had minimal impact on error metrics such as the CRPS. We defer a more comprehensive analysis of stochastic sampling to future work. Our algorithm can, in principle, be used to fully denoise multiple snapshots at once (e.g., output two snapshots per iteration when $N=\frac{1}{2}$). This could be useful in certain modalities or problems where subsequent snapshots are highly correlated, but is left to future work too.
@@ C.4 Temporal noise prior
In our training and sampling algorithms (Alg. 1-3), we use i.i.d. noise to corrupt the first window as well as append new pure-noise snapshots when sliding the window. As mentioned in the main text, it has been shown to be useful to sample temporally correlated noise snapshots. In our experiments, we use the simple “progressive noise model” technique proposed by ge2024noiseprior, which we found to improve results on the Navier-Stokes dataset compared to using i.i.d. noise (see Appendix˜E). This noise prior generates the noise for each snapshot autoregressively by combining a perturbed version of the previous noise with a new i.i.d. noise sample. Mathematically, the noise for the first snapshot is random, $\bm{\epsilon}^{(1)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, while for the following snapshots $k>1$:
@@
where $\alpha$ controls the correlation strength between snapshot noises and $\alpha=0$ corresponds to i.i.d. noise. In our experiments, we use $\alpha=1$ following the recommended practice in [ge2024noiseprior]. These temporally correlated noise samples replace the i.i.d. draws in lines 6 and 27 of Algorithm˜3. For example, line 6 becomes $\bar{\bm{x}}_{\rm{cur}}\leftarrow\hat{{\bm{y}}}_{1:W}^{\mathrm{init}}+\bm{\bar{\sigma}}(t_{\rm{cur}})\bm{\epsilon}_{1:W}$, where $\bm{\epsilon}_{1:W}:=(\bm{\epsilon}^{(1)},\dots,\bm{\epsilon}^{(W)})$. Note that the noise prior is applied during both training and sampling, enabling the model to learn how to deal with temporally correlated noise samples. A more comprehensive study of alternative noise priors in the context of ERDM is left to future work.
@@ D.1 Dataset details
We use the Navier-Stokes dataset with four obstacles introduced by otness21nnbenchmark. Its license is CC BY 4.0. We also use the ERA5 dataset provided by Weatherbench-2 [rasp2023weatherbench2]. Its license is CC BY 4.0.
@@
We use the $1.5^{\circ}$ resolution of the data (a $240\times 121$ grid) provided by [rasp2023weatherbench2], with 69 prognostic (input and output) variables: Temperature (t), geopotential (z), specific humidity (q), and the u and v components of wind (u,v) over 13 pressure levels as well as the four surface variables 2m temperature (2t), mean sea level pressure (mslp), and u and v components of winds at 10m (10u, 10v). Following Weatherbench-2, the 13 levels are $\{50,100,150,200,250,300,400,500,600,700,850,925,1000\}$. Numbers after any variable abbreviation refer to the level in hPa. For example, z500 refers to geopotential at 500 hPa. Besides these prognostic variables, we also use six static maps as additional inputs: A land-sea mask, soil type, orography, and $\sin(lat),\cos(lat)\cos(lon),\cos(lat)\sin(lon)$, where $lat,lon$ are the latitude and longitude of each location. We train our models on $12$-hourly data, as in [price2023gencast], from 1979 to 2020 and evaluate the models based on 64 initial conditions at 00/12 UTC, evenly spaced in 2021.
@@ D.2 Implementation details and hyperparameters
For all experiments and baselines that we train, we use the same basic ADM U-Net [dhariwal2021diffusion] architecture, only modified where necessary depending on the method. We always use $4$ down- and up-sampling blocks, with channel multipliers $1,2,3,4$. The base channel dimension is $64$ for Navier-Stokes and $256$ for ERA5 experiments. We use attention layers at the two lowest spatial resolutions. We use a dropout rate of 0.15 for Navier-Stokes, but disable it for ERA5. Other architectural hyperparameters are kept to the defaults from EDM. This results in a parameter count of $29.5$ million for Navier-Stokes, and $517$ million for ERA5. For the models operating on sequence, ERDM and EDM with $W>1$, we integrate causal temporal attention layers into the 2D U-Net before each down- and up-sampling block. This increases the respective parameter counts slightly to $31.5$ and $537$ million for Navier-Stokes and ERA5, respectively. For ERA5, we use circularly padded convolutions along the longitude dimension (and zero-padding for the latitude dimension) to better respect the spherical structure of the Earth.
@@
We train each model for 300 epochs with a global effective batch size of 32, leading to almost $57,000$ gradient steps. The effective batch size is kept the same by balancing per-GPU batch size and gradient accumulation steps appropriately over different training runs. We use a cosine learning rate schedule with a linear warmup period of 10 epochs and a peak learning rate of $6\times 10^{-4}$. We use the AdamW optimizer with a weight decay of $10^{-4}$. During evaluation, we use an exponential moving average (EMA)–based on an EMA decay rate of 0.9950.995–version of the model. Because the input spatial grid, $221\times 42$, after four halvings does not result in an even number–a problem for the residual connections in the upsampling blocks–we bilinearly interpolate it to a $256\times 64$ grid at the input layer (and back before the final convolution layer).